# Towards a Real-World Aligned Benchmark for Unlearning in Recommender Systems

### v1

@@ 3. Preliminary Results
alcohol 25% SCIF clip 0.2162 0.3170 0.2210 0.2518 0.7152 0.8282 95.01 0.00% 0.00% 0.6006
alcohol 50% SCIF clip 0.2135 0.3144 0.2197 0.2509 0.7069 0.8226 95.01 0.00% 0.00% 0.7062
alcohol 75% SCIF clip 0.2128 0.3143 0.2206 0.2520 0.7055 0.8235 95.01 0.00% 0.00% 0.8098
alcohol 100% SCIF clip 0.2133 0.3141 0.2220 0.2529 0.7052 0.8222 95.01 0.00% 0.00% 0.8810
@@
meat 25% SCIF clip 0.2176 0.3170 0.2220 0.2521 0.7188 0.8286 95.36 0.00% 0.00% 0.6509
meat 50% SCIF clip 0.2147 0.3166 0.2189 0.2505 0.7099 0.8249 95.36 0.00% 0.00% 0.7082
meat 75% SCIF clip 0.2142 0.3149 0.2189 0.2503 0.7078 0.8235 95.36 0.00% 0.42% 0.7154
meat 100% SCIF clip 0.2130 0.3142 0.2182 0.2497 0.7053 0.8221 95.36 0.00% 0.16% 0.7664
@@
Preliminary results for this can be seen in rows with Algorithm SCIF clip. Due to time constraints these results only stem from using one seed instead of averaging over five seeds, and $c=1$ is chosen heuristically. Further experiments need to be done to find the best value of $c$ and to confirm the results for this seed. SCIF clip outperforms the other algorithms in model utility and unlearning completeness metrics. However, SCIF clip is an order of magnitude slower than Fanchuan and Fanchuan might perform better than SCIF clip given more time.
The general unlearning algorithms Fanchuan and Kookmin perform well. They achieve utility levels comparable to the retrained models, with a similarly effective reduction in sensitive item predictions. Kookmin performs slightly better in most cases. Furthermore, we observe that the average time to unlearn the interactions of a single user is in the range of six seconds for Fanchuan. From the KL divergence over the forget set, we can see that the predictions of Kookmin are always closer in distribution to the retrained model than the predictions from Fanchuan. In summary, this experiment provides first evidence that our research questions from Section 2.2 can be positively answered with a new benchmark. The experiment indicates that unlearning also works for NBR models (RQ1) with many small unlearning requests (RQ3). We observe a case where a custom-designed unlearning algorithm for recommender systems outperforms general unlearning algorithms (RQ2). This only holds when applying update step norm clipping for the custom-designed unlearning algorithm. Without this modification, the performance drops significantly. Furthermore, unlearning algorithms can be executed with a low latency of several seconds only, which demonstrates the practical potential for deploying unlearning in real-world production systems (RQ4).

### v2

@@ 3. Preliminary Results
alcohol 25% SCIF clip 0.2175 0.3173 0.2225 0.2528 0.6879 0.7888 95.67 0.00% 0.37% 0.6082
alcohol 50% SCIF clip 0.2145 0.3152 0.2208 0.2518 0.6767 0.7851 95.67 0.00% 0.81% 0.7289
alcohol 75% SCIF clip 0.2142 0.3145 0.2222 0.2528 0.6751 0.7840 95.67 0.05% 0.71% 0.8347
alcohol 100% SCIF clip 0.2148 0.3146 0.2244 0.2546 0.6754 0.7840 95.67 0.00% 0.44% 0.8922
@@
meat 25% SCIF clip 0.2189 0.3182 0.2240 0.2539 0.6879 0.7902 96.36 0.12% 2.91% 0.5942
meat 50% SCIF clip 0.2163 0.3169 0.2213 0.2521 0.6813 0.7868 96.36 0.12% 1.95% 0.6594
meat 75% SCIF clip 0.2151 0.3157 0.2210 0.2520 0.6780 0.7854 96.36 0.08% 1.54% 0.7054
meat 100% SCIF clip 0.2145 0.3151 0.2213 0.2521 0.6760 0.7841 96.36 0.03% 1.30% 0.7466
@@
Results and discussion. We list the results in Table 2. The performance of the adapted unlearning method SCIF is volatile. For some seeds the model performance collapsed, and for almost all users unlearning items were predicted. We attribute this observation to misguided update steps originating from an ill-conditioned Hessian approximation, which is confirmed by the large $\ell_{2}$-norm of the parameter update. Large update steps are rare but harm all subsequent updates due to the incremental nature of sequential unlearning. With one seed, SCIF predicted 0% items for the category meat, but the model utility was worse compared to other approaches. We measured the $\ell_{2}$-norm of the parameter update steps. In each training run, there were some update steps having norm >100>100 while the average norm over a run was close to $1$. A reason for that can be an ill-conditioned Hessian. Approaches to circumvent the big update steps are clipping update steps by norm or a bigger damping for the Hessian vector product calculation when update steps get too big. We decided to implement norm clipping to a maximum of $1$ for the update steps. Let $\Delta\tilde{\theta}:=-H_{\tilde{\theta}}^{-1}\nabla_{\tilde{\theta}}\left(\frac{1}{2+bs}\left(-\ell(z,\theta)+\ell(\bar{z},\theta)+\sum_{i=1}^{bs}\ell(z_{i},\theta)\right)\right)$. We clip the update step norm to a maximum of $c\geq 0$:
@@
where $||\cdot||_{2}$ is the $\ell_{2}$-norm of the input tensor flattened into a 1-dimensional vector.
Preliminary results for this can be seen in rows with Algorithm SCIF clip. Due to time constraints $c=1$ is chosen heuristically. Further experiments need to be done to find the best value of $c$ and to confirm the results for this seed. SCIF clip outperforms the other algorithms in model utility and unlearning completeness metrics. However, SCIF clip is an order of magnitude slower than Fanchuan and Fanchuan might perform better than SCIF clip given more time.
The general unlearning algorithms Fanchuan and Kookmin perform well. They achieve utility levels comparable to the retrained models, with a similarly effective reduction in sensitive item predictions. Kookmin performs slightly better in most cases. Furthermore, we observe that the average time to unlearn the interactions of a single user is in the range of six seconds for Fanchuan. From the KL divergence over the forget set, we can see that the predictions of Kookmin are always closer in distribution to the retrained model than the predictions from Fanchuan.
In summary, this experiment provides first evidence that our research questions from Section 2.2 can be positively answered with a new benchmark. The experiment indicates that unlearning also works for NBR models (RQ1) with many small unlearning requests (RQ3). We observe a case where a custom-designed unlearning algorithm for recommender systems outperforms general unlearning algorithms (RQ2). This only holds when applying update step norm clipping for the custom-designed unlearning algorithm. Without this modification, the performance drops significantly. Furthermore, unlearning algorithms can be executed with a low latency of several seconds only, which demonstrates the practical potential for deploying unlearning in real-world production systems (RQ4).